OpenStudy (anonymous):

Write the equation of an exponential function that has a horizontal asymptote at y=3, and a y-intercept at (0,4) and passes through the point (-1,6). Please explain how to go through this, i dont want only answers

OpenStudy (sirm3d):

 \[\large y=a\cdot b^{-x}+c\] the line \(y=c\) is the asymptote

OpenStudy (anonymous):

i confused on how i would get a or b, i have this so far: +3 lol and nothing else

OpenStudy (sirm3d):

when x=0, y=4 \[4=a\cdot b^0+3\] but b^0=1

OpenStudy (sirm3d):

so a=1 finally, when x=-1, y=6 solve b in \[6=b^{-1}+3\]

OpenStudy (sirm3d):

uhm, i mean \[6=b^{-(-1)}+3\] or \[6=b+3\]

OpenStudy (anonymous):

okay so b is 3 y=3^-x +3?